# Euclid preparation

### v1

@@ 4.1 Optical quality of RGS000, RGS180 & BGS000
Because the asymmetric-Gaussian profile may not be able to catch the faint extended tail of the PSF, we did not use an analytical estimate of EE80, as we used for the EE50, but instead evaluated the EE80 after having subtracted the contributions from the fiber-core and background in the image, both derived from the model fitting. In this case EE80 is estimated by searching for the radius of the circle – centred onto the PSF centroid position – that encapsulates 80% of the total signal in the background-free image. However, when working with the two other models, to avoid introducing errors when subtracting the background or fibre-core model from the PSF image, the EE50 and EE80 were instead evaluated on the fitted model. In those cases, the radii were determined by evaluating the apertures containing 50% and 80% of the total signal from the PSF model, using either the dual-Gaussian or Moffat components of the model, and excluding both background and fibre-core contributions. However, as the monochromator did not have infinitely narrow bandwidth, and hence PSFs are slightly extended along the dispersion direction, both the EE50 and EE80 are slightly overestimated as this estimation disregards PSF asymmetry.
@@
The reduced -distributions of the three models have a median value of 1.04 for the asymmetric-Gaussian model, 0.97 for the dual-Gaussian model, and 1.02 for the Moffat profile. These values are too similar to prefer one model over the others. For every model a few cases show values larger than 1000. These were found to be correlated between the three models and are associated with the PSF being located close to hot or pixels with ‘defects’ like cosmic ray hits, Random Telegraph Noise, etc… that were not properly accounted for in our data reduction. We subsequently excluded these cases with biased fit results, removing PSFs with reduced which result in excluding about 5% of the PSFs. Additionally, we rejected any PSF with an unphysically large maximum signal 10 000 e- s – given the photon flux provided by the optical ground equipment setup.
@@
Figure 5 shows the relative differences in the measured position of the PSF between the different PSF models. With a standard deviation lower than 0.1 pixels and a mean of (7–10) 10-3 pixels, depending on the histogram, we concluded that all three models predict the same PSF centroid position with an error on the order of one tenth of a pixel. As those three models locate PSF at roughly the same position, it is safe to consider they all identify the same PSF which allowes a comparison of the three profiles.
@@
Despite the differences in the profile definition, the three models provide similar EE values. Given the goal to assess the NISP spectroscopy image quality from PSF size we conclude that these models overall provide robust estimates.
@@
In Fig. 6 the overall image quality is shown, expressed as mean and spread of PSF EE50 and EE80 across the NISP FoV, derived from the three PSF models. These are compared to an ideal optical diffraction-limited system (shaded area) and to the scientific requirements (dashed area). The boxes for each model represent mean, and first and third quartile of their values across the FoV – they are the combined result of the accuracy of our measurement and the actual PSF-size variations. Similarly, the width of the ideal model band shows the variations of the theoretical PSF across the FoV. One can see from this figure that the image quality of the NISP grisms is close to the theoretical expectation from an ideal system, with EE50 0.6 times smaller than the scientific requirements at 1500 nm. A more detailed investigation of the four target fields shows variations in the median value of both EE50 and EE80, suggesting a smooth variation of PSF size across the FoV. However, these are smaller than the spread due to measurement uncertainties and are being disregarded.
@@
One has to keep in mind that the data presented here were taken without the Euclid telescope system, and hence the telescope PSF has to be accounted for when extrapolating the measured NISP performance to in-flight conditions. Ground test of the Euclid payload module, which contains both instruments as well as the whole Euclid optical system, were conducted in summer 2022 at ‘Centre Spatial de Liège’ (CSL) by Airbus Defense and Space. Although these tests are outside the scope of this paper, we can report that while Euclid’s telescope was prone to gravity which stressed the mechanical structure of the primary mirror, analysis did not revealed any degradation beyond expectations.
5 Addressing RGS270 Non-Conformity
@@
However, during our ground test campaign at instrument level we observed that images from grism RGS270 were partially defocused. This can be seen in Fig. 9 which compares Fabry–Pérot spectrograms acquired with RGS000 and RGS270 during NISP ground test campaign. Focusing on the middle image presenting the RGS270 2D spectrogram, one can see that the bluest part of the spectrum (left-hand side) has a narrow and well focused PSF. In contrast to this, with increasing wavelength towards the right, the PSF gets successively wider and at the reddest end even shows a ring-like appearance. There is clearly an increasing defocus with increasing wavelength.
@@
To obtain the additional spectral orientation angles needed for decontamination of overlapping spectra, the survey solution involved to modify the survey strategy using NISP settings that rotate the NI-GWA by away from its nominal positioning of the grisms with respect to the centre of the beam, when observing either with RGS000 or RGS180. This in turn creates spectra tilted by vs. the nominal orientation of spectra from RGS000 and RGS180. After an in-depth simulation of different survey strategies we ended up with an optimal spectroscopic observation sequence. Each field is observed four times, with grisms used in the following order: RGS000 RGS180+ RGS000– RGS180.
The advantage of this observation sequence, executed in the ‘K-pattern’ of telescope dithering offsets for each sky position (Euclid Collaboration: Scaramella et al., 2022; Euclid Collaboration: Mellier et al., 2025), creates a geometry formed by the stacked spectra that offers even more dispersion angles than the initial observation sequence, for disentangling spectra during the offline data processing on ground. The largest drawback is that by slight rotation angle of the grism wheel, the nominal aperture of the grism – defined by a baffle installed in front of the grating – is shifted away from the nominal beam centre, introducing slight vignetting at the edge of the FoV. This vignetting has been estimated from Zemax simulation to result in 10% flux losses at field angle, i.e. distance from the centre of the FoV, which is reduced to flux losses at field angle . On-sky simulation, involving the official Euclid data reduction pipeline, demonstrated that this vignetting level is largely compensated by the detector quantum efficiency and an optical throughput which are both higher than scientifically required, and that the spatial structure can also easily be accounted for in the pipeline. Larger tilts were also investigated but the increasing vignetting level quickly reduces the options for modifying the Euclid survey strategy along this line.
@@ 6.1 On-ground spectral dispersion calibration
To estimate spectral dispersion functions of the NISP grisms, the Fabry–Pérot etalon spectrum was projected with the telescope simulator to 144 individual points on the NISP FoV. For every pointing a short exposure of 13 s was taken before the telescope simulator moved to the next field point. Additionally, at all positions corresponding to the centre of a NISP detector, we also used the Argon spectral lamp to obtain spectra as reference and control.
@@
Before going into the details of the calibration, Fig. 10 present examples of the -order spectrogram for the blue and red grisms acquired during the NISP ground tests. The middle panels each show and compare spectrograms of the Fabry–Pérot etalon with those of the Argon spectral lamp. The top panels show reference Argon spectra in vacuum in these wavelength ranges, taken from the NIST database (Kramida et al., 2022). By comparing the reference Argon spectra with those observed with NISP (middle and bottom panels) one can identify the strongest Argon emission lines present in both the and bandpasses. The middle panels also identify the Fabry–Pérot transmission peaks. For clarity, we limit the labelling to two peaks but based on the Fabry–Pérot calibration data all other transmission peaks were clearly identified222For reference see the calibrated Fabry–Pérot Spectral Energy Distribution (SED) in Fig. 2. The middle and bottom panels do not cover the spectrogram of the -order which is located about 500 pixel away off the left-hand side.
One note, upon close inspection of the bottom panels, that both show a slight periodic excess. This is a charge persistence signal from the previous Fabry–Pérot etalon exposure. However, persistence did not affect the exposure of the Fabry–Pérot etalon itself, since the pointing position was changed between subsequent exposures and the Fabry–Pérot light source was brighter than the spectral lamp.
@@
The coordinates of the Fabry–Pérot transmission peaks were measured by fitting the observed peaks with the PSF profile described in Sect. 4, providing position with an accuracy of of a pixel.
@@
The transmission of the optical ground equipment has some residual knowledge uncertainties. To limit the impact of this, the Fabry–Pérot template is constructed by taking the averaged profile of its -order spectrogram measured with the corresponding grism, corrected by the -order transmission efficiency. In this way the template is considered to be representative of the photon flux at the entrance pupil of the NISP instrument. To account for the -order, the template is then multiplied by the -order transmission efficiency (see Fig. 3) and convolved with a PSF profile before being adjusted to the 2D images of the -order. Parameters of the fit are the centroid position of the template, defined to be the central wavelength of the grism band-pass, where the -order transmission efficiency is at its minimum, a stretch factor converting wavelength to pixel, the dispersion direction angle, and parameters of the PSF profile defined in Eq. 1 which are accounting for the contribution of the fibre core.
@@
An example of such a template fit is shown in Fig. 11 which presents the measured image of the Fabry–Pérot -order spectrogram in (A) with the model in (B) and the residual in (C) and (D). Due to the prisms onto the NISP’s grisms, the -order are slightly dispersed, extending over several tens of pixels. Additionally, the grating’s blaze function is optimized to maximize transmission in the -order – around for the red grisms and for the blue grism – while minimizing transmission at the same wavelengths for the -order. As a result, instead of appearing as a point-like feature, as would be expected for a purely dispersive grating, NISP’s -order takes the form of an elongated, double-peaked structure. We tested the different PSF profiles we listed in Sect. 4 and concluded that they all provide similar results without any of them outperforming the others. As for modelling the monochromatic PSF, the usage of either of the profiles described in Sect. 4 led to similar centroiding of the -order with negligible impact on the derived calibration parameters. The -order position errors at were determined using -profiling and subsequently propagated to the centroid position of the Fabry–Pérot transmission peaks during the calibration of individual -order spectrograms. Analyzing the distribution of measured uncertainties in the -order positions, we obtain mean uncertainties of with a standard deviation of pixels in the cross-dispersion direction and with a standard deviation of pixels in the dispersion direction. This highlights that positioning is more precise in the cross-dispersion direction, as the profile of the -order is significantly thinner in this direction, leading to better constraint.
@@
Coefficients of the Chebyshev polynomial as described in Eq. 2 are evaluated from a recursive -fit. In the recursion, outliers were rejected at every iteration, each identified as spectral features located 5 away from the fitted spectral trace. Recursion stopped when no further outlier were identified. This allowed us to automatically reject some of the Fabry–Pérot transmission peaks that were insufficiently characterized by our PSF modelling due to nearby hot or bad pixels. Chebyshev polynomials where expanded up to the third order in the cross-dispersion direction and up to the fourth order in the dispersion direction. For the blue grism calibration the expansion was limited to the third order in both directions. The level of the expansion order was chosen in order for the averaged reduced to be the closest to . With the above expansion order, the averaged reduced is of the order of . However, we noticed that the distribution is skewed toward lower values. This is due to the limited accuracy of the template fit on a few of the -order images which propagated to the -order spectra and leads to large uncertainties in the relative position errors of some peaks’ positions, .
@@
Up to this point, the Fabry–Pérot spectrograms were modelled independently of each other, leading to calibration coefficients that smoothly vary across the FoV depending on the location of the order. To complete the calibration, the spatial dependency of each of the coefficients is then fitted by a bi-dimensional Chebyshev polynomial defined as
@@
Here are the limits of the focal plane array, set to for both the and axes. Again, the evaluation of the parameters is carried out by a fit. The tables that summarise the calibration coefficients for each of the grisms are presented in Appendix A.
Theoretically, in a perfect optical system with constant grating groove, one expects the spectrogram to rotate around their un-deflected wavelength when the grating is rotated around its centre. However, with the complexity of the NISP instrument itself, the mechanical uncertainties on the wheel position after its rotation, and on the pointing of the telescope simulators, it was not possible to precisely measure and confirm the position of the rotation centre for each of the spectrograms. This choice was made to minimize the bias induced by the uncertainty on the location of each spectrum’s centre of rotation. We tested the hypothesis that the rotation centre of the spectrograms was either their undeflected wavelength or the -order333Which, strictly speaking, is not physically possible given the grism design, but this was nevertheless also tested, since at first approximation the -order could be a proxy for the rotation centre.. However, the calibration error derived from such a nominal position was found to be outside of our scientific requirement once it was applied to the rotated grisms, despite being calculated at sub-pixel level.
@@
Additionally, we noticed that every spectrum for which the and orders were located on two different detectors shows a local offset in the estimated values of both the and coefficients, while the higher order coefficients of the parametrisation remained compatible with the coefficients estimated for any spectrogram fully falling onto a single detector. This offset behaves as if those spectra have a longer separation between the position of their and orders, breaking the otherwise rather smooth variation of the and coefficients across the field of view. Instead, we concluded that this offset has its origin in a bias induced by a limited precision of the detector metrology: this provides physical positions of the NISP detectors within the focal plane and was used to convert position measured in pixels on a given detector to the position in the NISP focal plane reference frame. However, the detector metrology was only determined at room temperature by measuring the position of reference marks engraved on the mechanical structure of each detector, using an optical camera mounted on metrology bench. This method only provided a precision of the order of one pixel (i.e. ). The NISP focal plane metrology at operational temperature was therefore computed relying on the thermal model of the NISP instrument to predict, with limited precision, how thermal stress would modify the focal plane.
We did try to re-calibrate the detector metrology at cold using the spectroscopic data themself by aligning -order spectra dispersed over multiple detectors. However, the small number of sources available (i.e. one single spectrum per NISP exposure in the entire NISP focal plane) did not allow a re-computation with a sufficient precision for the spectroscopic calibration. We therefore left our cold reference-metrology as is and checked that the calibration parameters remain within 3 when computing calibration parameters with or without the spectra showing an offset in the and coefficients.

### v2

@@ 4.1 Optical quality of RGS000, RGS180 & BGS000
Because the asymmetric-Gaussian profile may not be able to catch the faint extended tail of the point spread function, we did not use an analytical estimate of EE80, as we used for the EE50, but instead evaluated the EE80 after having subtracted the contributions from the fibre-core and background in the image, both derived from the model fitting. In this case EE80 is estimated by searching for the radius of the circle – centred onto the point spread function centroid position – that encapsulates 80% of the total signal in the background-free image. However, when working with the two other models, to avoid introducing errors when subtracting the background or fibre-core model from the point spread function image, the EE50 and EE80 were instead evaluated on the fitted model. In those cases, the radii were determined by evaluating the apertures containing 50% and 80% of the total signal from the point spread function model, using either the dual-Gaussian or Moffat components of the model, and excluding both background and fibre-core contributions. However, as the monochromator did not have infinitely narrow bandwidth, and hence point spread functions are slightly extended along the dispersion direction, both the EE50 and EE80 are slightly overestimated as this estimation disregards point spread function asymmetry.
The fits of the point spread function using the models presented in Eq. 1 are performed on each individual point spread function exposure. The spatial sampling provided by dithering allows us to derive the statistical distribution of the point spread function model parameters, which will be discussed in the following paragraph.
@@
The reduced -distributions of the three models have a median value of 1.04 for the asymmetric-Gaussian model, 0.97 for the dual-Gaussian model, and 1.02 for the Moffat profile. These values are too similar to prefer one model over the others. For every model a few cases show values larger than 1000. These were found to be correlated between the three models and are associated with the point spread function being located close to hot or pixels with ‘defects’ like cosmic ray hits, Random Telegraph Noise, etc… that were not properly accounted for in our data reduction. We subsequently excluded these cases with biased fit results, removing point spread functions with reduced which result in excluding about 5% of the point spread functions. Additionally, we rejected any point spread function with an unphysically large maximum signal 10 000 e- s – given the photon flux provided by the optical ground equipment setup.
@@
Figure 5 shows the relative differences in the measured position of the point spread function between the different point spread function models. With a standard deviation lower than 0.1 pixels and a mean of (7–10) 10-3 pixels, depending on the histogram, we concluded that all three models predict the same point spread function centroid position with an error on the order of one tenth of a pixel. As those three models locate point spread function at roughly the same position, it is safe to consider they all identify the same point spread function which allows a comparison of the three profiles.
@@
Despite the differences in the profile definition, the three models provide similar encircled energy values. Given the goal to assess the NISP spectroscopy image quality from point spread function size we conclude that these models overall provide robust estimates.
@@
In Fig. 6 the overall image quality is shown, expressed as mean and spread of point spread function EE50 and EE80 across the NISP field-of-view, derived from the three point spread function models. These are compared to an ideal optical diffraction-limited system (shaded area) and to the scientific requirements (dashed area). The boxes for each model represent mean, and first and third quartile of their values across the field-of-view – they are the combined result of the accuracy of our measurement and the actual point spread function-size variations. Similarly, the width of the ideal model band shows the variations of the theoretical point spread function across the field-of-view. One can see from this figure that the image quality of the NISP grisms is close to the theoretical expectation from an ideal system, with EE50 0.6 times smaller than the scientific requirements at 1500 nm. A more detailed investigation of the four target fields shows variations in the median value of both EE50 and EE80, suggesting a smooth variation of point spread function size across the field-of-view. However, these are smaller than the spread due to measurement uncertainties and are being disregarded.
@@
One has to keep in mind that the data presented here were taken without the Euclid telescope system, and hence the telescope point spread function has to be accounted for when extrapolating the measured NISP performance to in-flight conditions. Ground test of the Euclid payload module, which contains both instruments as well as the whole Euclid optical system, were conducted in summer 2022 at ‘Centre Spatial de Liège’ by Airbus Defence and Space. Although these tests are outside the scope of this paper, we can report that while Euclid’s telescope was prone to gravity which stressed the mechanical structure of the primary mirror, analysis did not reveal any degradation beyond expectations.
4.2 Addressing RGS270 Non-Conformity
@@
However, during our ground test campaign at instrument level we observed that images from grism RGS270 were partially defocused. This can be seen in Fig. 9 which compares Fabry–Pérot spectrograms acquired with RGS000 and RGS270 during NISP ground test campaign. Focusing on the middle image presenting the RGS270 2D spectrogram, one can see that the bluest part of the spectrum (left-hand side) has a narrow and well focused point spread function. In contrast to this, with increasing wavelength towards the right, the point spread function gets successively wider and at the reddest end even shows a ring-like appearance. There is clearly an increasing defocus with increasing wavelength.
@@
To obtain the additional spectral orientation angles needed for decontamination of overlapping spectra, the survey solution involved to modify the survey strategy using NISP settings that rotate the grism wheel by away from its nominal positioning of the grisms with respect to the centre of the beam, when observing either with RGS000 or RGS180. This in turn creates spectra tilted by vs. the nominal orientation of spectra from RGS000 and RGS180. After an in-depth simulation of different survey strategies we ended up with an optimal spectroscopic observation sequence. Each field is observed four times, with grisms used in the following order: RGS000 RGS180+ RGS000– RGS180.
The advantage of this observation sequence, executed in the ‘K-pattern’ of telescope dithering offsets for each sky position (Euclid Collaboration: Scaramella et al., 2022; Euclid Collaboration: Mellier et al., 2025), creates a geometry formed by the stacked spectra that offers even more dispersion angles than the initial observation sequence, for disentangling spectra during the offline data processing on ground. The largest drawback is that by slight rotation angle of the grism wheel, the nominal aperture of the grism – defined by a baffle installed in front of the grating – is shifted away from the nominal beam centre, introducing slight vignetting at the edge of the field-of-view. This vignetting has been estimated from Zemax simulation to result in 10% flux losses at field angle, i.e. distance from the centre of the field-of-view, which is reduced to flux losses at field angle . On-sky simulation, involving the official Euclid data reduction pipeline, demonstrated that this vignetting level is largely compensated by the detector quantum efficiency and an optical throughput which are both higher than scientifically required, and that the spatial structure can also easily be accounted for in the pipeline. Larger tilts were also investigated but the increasing vignetting level quickly reduces the options for modifying the Euclid survey strategy along this line.
@@ 5.1 On-ground spectral dispersion calibration
To estimate spectral dispersion functions of the NISP grisms, the Fabry–Pérot etalon spectrum was projected with the telescope simulator to 144 individual points on the NISP field-of-view. For every pointing a short exposure of 13 s was taken before the telescope simulator moved to the next field point. Additionally, at all positions corresponding to the centre of a NISP detector, we also used the Argon spectral lamp to obtain spectra as reference and control.
@@
Before going into the details of the calibration, Fig. 10 presents examples of the -order spectrogram for the blue and red grisms acquired during the NISP ground tests. The middle panels each show and compare spectrograms of the Fabry–Pérot etalon with those of the Argon spectral lamp. The top panels show reference Argon spectra in vacuum in these wavelength ranges, taken from the NIST database (Kramida et al., 2022). By comparing the reference Argon spectra with those observed with NISP (middle and bottom panels) one can identify the strongest Argon emission lines present in both the and bandpasses. The middle panels also identify the Fabry–Pérot transmission peaks. For clarity, we limit the labelling to two peaks but based on the Fabry–Pérot calibration data all other transmission peaks were clearly identified222For reference see the calibrated Fabry–Pérot spectral energy distribution in Fig. 2. The middle and bottom panels do not cover the spectrogram of the -order which is located about 500 pixel away off the left-hand side.
One note that, upon close inspection of the bottom panels, that both show a slight periodic excess. This is a charge persistence signal from the previous Fabry–Pérot etalon exposure. However, persistence did not affect the exposure of the Fabry–Pérot etalon itself, since the pointing position was changed between subsequent exposures and the Fabry–Pérot light source was brighter than the spectral lamp.
@@
The coordinates of the Fabry–Pérot transmission peaks were measured by fitting the observed peaks with the point spread function profile described in Sect. 4, providing position with an accuracy of of a pixel.
@@
The transmission of the optical ground equipment has some residual knowledge uncertainties. To limit the impact of this, the Fabry–Pérot template is constructed by taking the averaged profile of its -order spectrogram measured with the corresponding grism, corrected by the -order transmission efficiency. In this way the template is considered to be representative of the photon flux at the entrance pupil of the NISP instrument. To account for the -order, the template is then multiplied by the -order transmission efficiency (see Fig. 3) and convolved with a point spread function profile before being adjusted to the 2D images of the -order. Parameters of the fit are the centroid position of the template, defined to be the central wavelength of the grism band-pass, where the -order transmission efficiency is at its minimum, a stretch factor converting wavelength to pixel, the dispersion direction angle, and parameters of the point spread function profile defined in Eq. 1 which are accounting for the contribution of the fibre core.
@@
An example of such a template fit is shown in Fig. 11 which presents the measured image of the Fabry–Pérot -order spectrogram in (A) with the model in (B) and the residual in (C) and (D). Due to the prisms onto the NISP’s grisms, the -order are slightly dispersed, extending over several tens of pixels. Additionally, the grating’s blaze function is optimised to maximise transmission in the -order – around for the red grisms and for the blue grism – while minimising transmission at the same wavelengths for the -order. As a result, instead of appearing as a point-like feature, as would be expected for a purely dispersive grating, NISP’s -order takes the form of an elongated, double-peaked structure. We tested the different point spread function profiles we listed in Sect. 4 and concluded that they all provide similar results without any of them outperforming the others. As for modelling the monochromatic point spread function, the usage of either of the profiles described in Sect. 4 led to similar centroiding of the -order with negligible impact on the derived calibration parameters. The -order position errors at were determined using -profiling and subsequently propagated to the centroid position of the Fabry–Pérot transmission peaks during the calibration of individual -order spectrograms. Analyzing the distribution of measured uncertainties in the -order positions, we obtain mean uncertainties of with a standard deviation of pixels in the cross-dispersion direction and with a standard deviation of pixels in the dispersion direction. This highlights that positioning is more precise in the cross-dispersion direction, as the profile of the -order is significantly thinner in this direction, leading to better constraint.
@@
Coefficients of the Chebyshev polynomial as described in Eq. 2 are evaluated from a recursive -fit. In the recursion, outliers were rejected at every iteration, each identified as spectral features located 5 away from the fitted spectral trace. Recursion stopped when no further outlier were identified. This allowed us to automatically reject some of the Fabry–Pérot transmission peaks that were insufficiently characterized by our point spread function modelling due to nearby hot or bad pixels. Chebyshev polynomials were expanded up to the third order in the cross-dispersion direction and up to the fourth order in the dispersion direction. For the blue grism calibration the expansion was limited to the third order in both directions. The level of the expansion order was chosen in order for the averaged reduced to be the closest to . With the above expansion order, the averaged reduced is of the order of . However, we noticed that the distribution is skewed towards lower values. This is due to the limited accuracy of the template fit on a few of the -order images which propagated to the -order spectra and leads to large uncertainties in the relative position errors of some peaks’ positions, .
@@
Up to this point, the Fabry–Pérot spectrograms were modelled independently of each other, leading to calibration coefficients that smoothly vary across the field-of-view depending on the location of the order. To complete the calibration, the spatial dependency of each of the coefficients is then fitted by a bi-dimensional Chebyshev polynomial defined as
@@
Here are the limits of the focal plane array, set to for both the and axes. Again, the evaluation of the parameters is carried out by a fit. The tables that summarise the calibration coefficients for each of the grisms are presented in Appendix A. The advantage of Eq. 2 is that, through the set of coefficients , it simultaneously describes the spectral dispersion and the curvature of the spectra, including non-linear effects induced by optical field distortion. Furthermore, by modelling each coefficient over the entire field of view using Eq. 3, one directly obtains a two-dimensional mapping of both dispersion and curvature through the higher-order coefficients or greater than or equal to 2.
Theoretically, in a perfect optical system with constant grating groove, one expects the spectrograms to rotate around their undeflected wavelength when the grating is rotated around its centre. However, with the complexity of the NISP instrument itself, the mechanical uncertainties on the wheel position after its rotation, and on the pointing of the telescope simulators, it was not possible to precisely measure and confirm the position of the rotation centre for each of the spectrograms. The choice was made to minimise the bias induced by the uncertainty on the location of each spectrum’s centre of rotation by considering each grism’s orientation as an independent grism to be calibrated. We tested the hypothesis that the rotation centre of the spectrograms was either their undeflected wavelength or the -order333Which, strictly speaking, is not physically possible given the grism design, but this was nevertheless also tested, since at first approximation the -order could be a proxy for the rotation centre.. However, the calibration error derived from such a nominal position was found to be outside of our scientific requirement once it was applied to the rotated grisms, despite being calculated at sub-pixel level.
@@
Additionally, we noticed that every spectrum for which the and orders were located on two different detectors shows a local offset in the estimated values of both the and coefficients, while the higher order coefficients of the parametrisation remained compatible with the coefficients estimated for any spectrogram fully falling onto a single detector. This offset behaves as if those spectra have a longer separation between the position of their and orders, breaking the otherwise rather smooth variation of the and coefficients across the field of view. Instead, we concluded that this offset has its origin in a bias induced by a limited precision of the detector metrology: this provides physical positions of the NISP detectors within the focal plane and was used to convert positions measured in pixels on a given detector to positions in the NISP focal plane reference frame. However, the detector metrology was only determined at room temperature by measuring the position of reference marks engraved on the mechanical structure of each detector, using an optical camera mounted on a metrology bench. This method only provided a precision of the order of one pixel (i.e. ). The NISP focal plane metrology at operational temperature was therefore computed relying on the thermal model of the NISP instrument to predict, with limited precision, how thermal stress would modify the focal plane.
We try to re-calibrate the detector metrology at cold using the spectroscopic data themselves by aligning -order spectra dispersed over multiple detectors. However, the small number of sources available (i.e. one single spectrum per NISP exposure in the entire NISP focal plane) did not allow a re-computation with sufficient precision for the spectroscopic calibration. We therefore left our cold reference-metrology as is and checked that the calibration parameters remain within 3 when computing calibration parameters with or without the spectra showing an offset in the and coefficients.